Monthly Construction and Environmental Update

West End (SR 79)

  • Utility delay regarding repositioning of AT&T fiber optic lines. Utilities working on resolution but effect on project progress not known at this time.
    • Gas utility relocation work delayed due to AT&T fiber optic issue.
  • Jefferson County water line installation (under separate NPDES permit) on-going near Green Drive. BMPs installed by Jefferson County contractor.
  • Cross drain pipes and culvert extensions on south end of SR 79 have been completed.
  • Slope clearing/roadway widening west of SR79. BMPs in place.
  • Temporary Sediment Basins 31 and 32 (with pumps and skimmers) functioning properly.

Mainline and East End (Station 2150 to Station 2235 at SR 75)   

  • Inlets, ditch, and slope work on local-traffic access road near SR 75 completed.
  • Archaeology excavation at Station 2205 by rock overhang ongoing.
  • Soil Nail Wall construction remains ongoing.
  • Retaining Wall foundation work near Self Creek completed.
  • Retaining Wall precast panel erection underway.
  • Clearing from Station 2150+00 to Station 2180+00 completed.
  • Blasting of rock throughout mainline section performed as needed.
  • Perimeter and other appropriate BMPs installed.
  • Temporary, permanent, and post-construction BMPs remain in place.
    • Construction entrances remain stabilized.
    • Disturbed areas are temporarily stabilized with blown or hydraulically-applied mulch; diversion berms, slope drains, and sump excavations have been installed where needed.
    • Sediment barriers installed at buffers, stream and construction limits.
    • Six permanent and one temporary sedimentation basins with dewatering devices and baffles are functional.
    • Six Level spreaders associated with permanent sedimentation basins are functional.
    • Temporary slope drains, diversion berms, swales, sumps, wattles, ditch checks, and temporary seed/mulch installed along haul roads.

Environmental Status

  • 4 Rain Days: 10/3 (0.99”), 10/12 (0.30”), 10/14 (2.91”), 10/29 (0.63”).
  • Rain event on 10/14 exceeded the required design storm (2 yr-24 hour) for BMPs. ADEM was notified and damaged BMPs were repaired within 5 days. Turbid discharge was observed and reported to ADEM.
  • Total rainfall for October was approximately 4.83 inches